OpenStudy (anonymous):

(4 sqrt x^6)(12 sqrt^9)

OpenStudy (solomonzelman):

well... that's simplify the roots. \[\sqrt{x^{6}}=\sqrt{(x^{3})^2}=x^{3}\]\[\sqrt{x^{9}}=\sqrt{x \times (x^{4})^2}=\sqrt{x} \times \sqrt{(x^{4})^2}=\sqrt{x} \times x^{4}=x^4 \sqrt{x}\]

OpenStudy (solomonzelman):

So the problem you really have, is not \[(4 \sqrt{x^6} )~~(12 \sqrt{x^9})\]rather, (as we simplified the roots, it is... \[4 x^3 \times 12x^4\sqrt{x}\] Do you need help multiplying all of this, or not? Let me know if you do .

OpenStudy (anonymous):

so it would be x=9/4

OpenStudy (solomonzelman):

x is not going got be equal to anything. Weird ik, but it is not "x=" you are simplifying the product of the terms, that's all.

OpenStudy (solomonzelman):

\(\LARGE\color{blue}{ \bf 4x^{3} \times 12x^{4}\sqrt{x}= }\) \(\LARGE\color{blue}{ \bf 4\times 12 \times x^{3} \times x^{4} \times \sqrt{x}= }\) \(\LARGE\color{blue}{ \bf 48 \times x^{3} \times x^{4} \times \sqrt{x}= }\) \(\LARGE\color{blue}{ \bf 48 \times x^{7} \times \sqrt{x}= }\) \(\LARGE\color{blue}{ \bf 48x^{7}\sqrt{x}~. }\)

OpenStudy (anonymous):

in my original problem, its the square root to the 4th power and the square root to the 12 power. (not 4 and 12 times the power)

OpenStudy (solomonzelman):

CAn tyou write it out in here please? (like in numbers) ?